# Text-to-Audio Generation using Instruction-Guided Latent Diffusion Model

### v2

@@ 3.2. Preference Optimization for Language Models
Figure 1. An illustration of our pipeline for text-to-audio alignment. The top part depicts the preference dataset creation where three strategies are deployed to generate the undesirable audio outputs to the input prompts. These samples are further filtered to form Audio-alpaca. This preference dataset is finally used to align Tango using DPO-diffusion loss (Eq. 17), resulting Tango 2.
@@ 4.1.1. Audio Generation from Text Prompts
This strategy is aimed at prompts that describe some composition of sequence and simultaneity of events. To identify such prompts in AudioCaps’ training dataset, as a heuristics, we look for the following keywords in a prompt: before, after, then, or followed. We denote the set of such prompts as $\mathcal{T}_{2}$.
@@ 4.1.2. Ranking and Preference-Data Selection
Table 2. Statistic of Audio-alpaca
@@ 5.3. Evaluation Metrics
AudioLDM 2-Full-Large 2.33 1.58 8.14 0.44 3.54 3.16 5.82 1.09 6.60 0.49 3.65 3.41
Tango-full-FT 2.69 1.16 7.85 0.54 3.83 3.80 7.52 1.01 6.41 0.51 3.67 3.49
@@ 5.4. Main Results
5.5. Analyses
Table 6. GPT-4 prompt used to extract events or concepts from prompts.
@@ 5.5. Analyses
Example 1 - Input: An aircraft engine runs and vibrates, metal spinning and grinding occur, and the engine accelerates and fades into the distance Output: Firstly, an aircraft engine runs and vibrates. Then, I hear metal spinning and grinding. Then, the aircraft engine accelerates. Finally, the aircraft fades into the distance. So, here is the list of events that I observed: [”aircraft engine runs”, ”aircraft engine vibrates”, ”metal spinning”, ”metal grinding”, ”aircraft engine acclerates”, ”aircraft fades into the distance”]
Example 2 - Input: Bubbles gurgling and water spraying as a man speaks softly while crowd of people talk in the background Output: Firstly, I hear bubble gurgling. Also, I hear water spraying. Simultaneously, a man is speaking softly. Also, a crowd of people are talking in the background. So, here is the list of events that I observed: [”bubble gurgling”, ”water spraying”, ”a man is speaking softly”, ”crowd talking”]
Example 3 - Input: A man talking then meowing and hissing Output: Firstly, I hear a man talking. Subsequently, I hear meowing. I also hear hissing. So, here is the list of events that I observed: [”a man talking”, ”meowing”, ”hissing”]
@@
Results on Temporal and Multi-concept Prompts.
We analyze the performance of Tango and AudioLDM2 models in audio generation when text prompts contain multiple sequential events. For instance, Two gunshots followed by birds flying away then a boy laughing consists of three distinct events in a sequence. In contrast, “A man is snoring” lacks any temporal sequence. We partition the AudioCaps test set based on temporal identifiers such as ”while,” ”as,” ”before,” ”after,” ”then,” and ”followed,” creating two subsets: one with multiple sequential events and the other with no temporality. Our objective evaluation results for these subsets are shown in Table 5. Tango 2 achieves the highest scores of KL, CLAP, and IS on both temporal and non-temporal events, indicating a consistent trend of performance improvement. Similar trends are observed in subjective evaluations. We attribute this improvement to the addition of both temporal and non-temporal samples to Audio-alpaca.
Additionally, we categorize prompts based on the presence of multiple concepts or events, exemplified by phrases like A woman speaks while cooking. As underlined, this prompt contains two distinct events i.e., “sound of a woman speaking” and “sound of cooking”. Through manual scrutiny, we discovered that pinpointing prompts with such multi-concepts is challenging using basic parts-of-speech or named entity-based rules. Consequently, we rely on GPT-4 to extract various concepts or events from the prompts using in-context exemplars. The specific prompt is displayed in Table 6. To evaluate GPT-4’s performance on this task, we randomly selected outputs for 30 unique prompts and manually verified GPT-4’s annotations. No errors attributable to GPT-4 were found. In general, Tango 2 outperforms AudioLDM2 and Tango across most objective and subjective metrics, following Table 4. We proceed to visualize the CLAP scores of the models in Figure 3. This visualization illustrates that Tango 2 consistently outperforms the baselines as the number of events or concepts per prompt increases. In particular, Specifically, Tango closely matches the performance of Tango 2 only when the textual prompt contains a single concept. However, the disparity between these two models widens as the complexity of the prompt increases with multiple concepts.
@@
We conducted an ablation study on Audio-alpaca to delve into the impact of different negative data construction strategies. As depicted in Table 3, excluding the data samples generated by Strategies 2 and 3 notably diminishes the performance of Tango 2. This underscores the significance of event and temporal prompt perturbations.

### v3

@@ 3.2. Preference Optimization for Language Models
Figure 1. An illustration of our pipeline for text-to-audio alignment. The top part depicts the preference dataset creation where three strategies are deployed to generate the undesirable audio outputs to the input prompts. These samples are further filtered to form Audio-alpaca. This preference dataset is finally used to align Tango using DPO-diffusion loss (Eq. 17), resulting in Tango 2.
@@ 4.1.1. Audio Generation from Text Prompts
This strategy is aimed at prompts that describe some composition of sequence and simultaneity of events. To identify such prompts in AudioCaps’ training dataset, as a heuristics, we look for the following keywords in a prompt: while, before, after, then, or followed. We denote the set of such prompts as $\mathcal{T}_{2}$.
@@ 4.1.2. Ranking and Preference-Data Selection
Table 2. Statistics of Audio-alpaca.
@@ 5.3. Evaluation Metrics
AudioLDM 2-Full-Large 2.33 1.58 8.14 0.44 3.54 3.16 5.82 1.09 4.60 0.49 3.65 3.41
Tango-full-FT 2.69 1.16 7.85 0.54 3.83 3.80 7.52 1.01 4.87 0.51 3.67 3.49
@@ 5.4. Main Results
Table 6. GPT-4 prompt used to extract events or concepts from audio prompts.
@@
Example 1 -
Input: An aircraft engine runs and vibrates, metal spinning and grinding occur, and the engine accelerates and fades into the distance
Output: Firstly, an aircraft engine runs and vibrates. Then, I hear metal spinning and grinding. Then, the aircraft engine accelerates. Finally, the aircraft fades into the distance.
So, here is the list of events that I observed:
["aircraft engine runs", "aircraft engine vibrates", "metal spinning", "metal grinding", "aircraft engine acclerates", "aircraft fades into the distance"]
Example 2 -
Input: Bubbles gurgling and water spraying as a man speaks softly while crowd of people talk in the background
Output: Firstly, I hear bubble gurgling. Also, I hear water spraying. Simultaneously, a man is speaking softly. Also, a crowd of people are talking in the background.
So, here is the list of events that I observed:
["bubble gurgling", "water spraying", "a man is speaking softly", "crowd talking"]
Example 3 -
Input: A man talking then meowing and hissing
Output: Firstly, I hear a man talking. Subsequently, I hear meowing. I also hear hissing.
So, here is the list of events that I observed:
["a man talking", "meowing", "hissing"]
@@
Additionally, we categorize prompts based on the presence of multiple concepts or events, exemplified by phrases like A woman speaks while cooking. As underlined, this prompt contains two distinct events i.e., “sound of a woman speaking” and “sound of cooking”. Through manual scrutiny, we discovered that pinpointing prompts with such multi-concepts is challenging using basic parts-of-speech or named entity-based rules. Consequently, we task GPT-4 with extracting the various concepts or events from the prompts using in-context exemplars. The specific prompt is displayed in Table 6. To evaluate GPT-4’s performance on this task, we randomly selected 30 unique prompts and manually verified their annotations from GPT-4’s. No errors attributable to GPT-4 were found. In general, Tango 2 outperforms AudioLDM2 and Tango across most objective and subjective metrics, following Table 4. We proceed to visualize the CLAP scores of the models in Figure 3. This visualization illustrates that Tango 2 consistently outperforms the baselines as the number of events or concepts per prompt increases. In particular, specifically, Tango closely matches the performance of Tango 2 only when the textual prompt contains a single concept. However, the disparity between these two models widens as the complexity of the prompt increases with multiple concepts.
@@
We conducted an ablation study on Audio-alpaca to gauge the impact of different negative data construction strategies. As shown in Table 3, excluding the data samples from by strategies 2 and 3 notably diminishes the performance of Tango 2. This underscores the significance of event and temporal prompt perturbations.